# Uncertainty-Based Extensible Codebook for Discrete Federated Learning in Heterogeneous Data Silos

### v1

@@ 3.3 Loss Function
Since we introduce learnable codewords in our method, there are two parts of the loss function. For our task, we utilize cross-entropy as the loss function as follows:
$\mathcal{L}_{task}=-\sum_{class}y\log p$ (5)
where $p$ is the model output, and $y$ is the ground truth label.
For codebook optimization, akin to the strategy employed in VQ-VAE (Van Den Oord et al., 2017), we apply a stop gradient operation for the codeword update as follows:
@@
So, the total loss is:
$\mathcal{L}_{UEFL}=\mathcal{L}_{task}+\mathcal{L}_{code}$ (7)
@@ 3.4 Uncertainty Evaluation
As outlined in Section 3.1, evaluating model uncertainty is crucial for identifying data from heterogeneous distributions requiring supplementary codewords.
In our work, we utilize Monte Carlo Dropout (MC Dropout) (Gal & Ghahramani, 2016) for uncertainty evaluation, incorporating two dropout layers into our model for regularization purposes. Unlike traditional usage where dropout layers are disabled during inference to stabilize predictions, we activate these layers during testing to generate a variety of outcomes for uncertainty analysis. The inherent randomness of dropout facilitates the acquisition of diverse predictions, aiding in the assessment of uncertainty. This variability is quantified using predictive entropy, as described in Equation 8, which serves to measure the prediction dispersion across different evaluations effectively.
@@
A low predictive entropy value signifies model confidence, whereas a high value indicates increased uncertainty. In situations of high entropy, introducing new codewords and conducting additional training rounds are essential steps. Given the variability of uncertainty across datasets, establishing a fixed threshold is impractical. Instead, by analyzing all uncertainty values, we can benchmark against either the minimum or mean values to pinpoint target silos. Our experiments showed superior results when using the minimum value as a reference, thus guiding us to adopt the following threshold criterion:
@@ 4 Experimental Results
Dataset and Model Our UEFL is rigorously assessed using five datasets: MNIST, FMNIST, CIFAR10, GTSRB, and CIFAR100. To validate our framework’s robustness, we employ various augmentation techniques to engineer heterogeneous datasets. The original dataset forms the baseline data distribution, $\mathcal{D}_{1}$. For MNIST, augmentations include clockwise rotations of 50 degrees and counterclockwise rotations of 120 degrees, generating two novel data distributions, $\mathcal{D}_{2}$ and $\mathcal{D}_{3}$. For the remaining four datasets, Gaussian noise ($\mathcal{N}(0,10^{2})$) is added to $\mathcal{D}_{3}$ to enhance heterogeneity. Data silos for CIFAR100 contain 4000 images each, while the other datasets consist of 2000 images per silo.
For RGB datasets like GTSRB, CIFAR10, and CIFAR100, we adopt a pretrained VGG16 model. In contrast, for grayscale datasets such as MNIST and FMNIST, lacking pretrained models, we design a convolutional network comprising three ResNet blocks, training it from scratch. Initial codebook sizes are set to 32 for MNIST and 64 for the remaining datasets, with an equivalent number of codewords added in each subsequent iteration. While additional iterations may converge within 5 rounds, we extend this to 20 for enhanced experimental clarity. The uncertainty evaluation is conducted 20 times using a dropout rate of 0.1, with thresholds $\gamma$ set at 0.3 for MNIST, 0.1 for FMNIST, GTSRB, and CIFAR100, and 0.2 for CIFAR10, to fine-tune performance. These experiments are performed on a machine with two NVIDIA A6000 GPUs.
Evaluation Metrics For image classification, we utilize Top1 accuracy as the benchmark for both local and global models. To enable a straightforward comparison, particularly given the consistent performance across data silos within the same distribution, we calculate the mean accuracy (mA) across all silos for each distribution, as follows:
$mA=\frac{1}{M}\sum_{i=1}^{M}a_{i}$ (10)
where $M$ is the number of distributions, $a_{i}$ is the accuracy for the $i$th data silo from this distribution.
In addition, we evaluate entropy as model uncertainty as Equation 8. We also evaluate the perplexity (PPL) to show the utility of codewords as follows,
$PPL=exp(-\sum_{i=1}^{N}p_{i}\log p_{i})$ (11)
where $N$ is the number of codewords, and $p_{i}$ denotes the probability of the $i$th codeword occurring.
Similar to mA, we evaluate mean entropy (mE) as $mE=\frac{1}{M}\sum_{i=1}^{M}e_{i}$ and mean perplexity (mP) as $mP=\frac{1}{M}\sum_{i=1}^{M}PPL_{i}$,

### v4

@@ 3.3 Loss Function
Since we introduce learnable codewords in our method, there are two parts of the loss function. For our task, we utilize cross-entropy as the loss function. For codebook optimization, akin to the strategy employed in VQ-VAE (Van Den Oord et al., 2017), we apply a stop gradient operation for the codeword update as follows:
@@
The total loss $\mathcal{L}_{UEFL}$ is the summation of $\mathcal{L}_{task}$ and $\mathcal{L}_{code}$.
@@ 3.4 Uncertainty Evaluation
As outlined in Section 3.1, evaluating model uncertainty is crucial for identifying data from heterogeneous distributions requiring supplementary codewords. In our work, we utilize Monte Carlo Dropout (MC Dropout) (Gal & Ghahramani, 2016) for uncertainty evaluation, incorporating two dropout layers into our model for regularization purposes. Unlike traditional usage where dropout layers are disabled during inference to stabilize predictions, we activate these layers during testing to generate a variety of outcomes for uncertainty analysis. This variability is quantified using predictive entropy, as described in Equation 6, which serves to measure the prediction dispersion across different evaluations effectively.
@@
A low predictive entropy value signifies model confidence, whereas a high value indicates increased uncertainty. For high entropy, introducing new codewords and conducting additional training rounds are essential steps. Given the variability of uncertainty across datasets, establishing a fixed threshold is impractical. Instead, by analyzing all uncertainty values, we can benchmark against either the minimum or mean values to pinpoint target silos. Our experiments showed superior results when using the minimum value as a reference, thus guiding us to adopt the following threshold criterion:
@@
Uncertainty decreases consistently during training, making it an effective stopping criterion for codebook extension. Besides, we manually set the maximum number of iterations to 5. A comparison with Deep Ensembles (Lakshminarayanan et al., 2017) is provided in Appendix E.
Table 1: UEFL outperforms all baselines on heterogeneous data. DisTrans lacks a Dropout layer, rendering it incapable of evaluating uncertainty. More results are in Appendix C.
Methods MNIST FMNIST GTSRB CIFAR10 CIFAR100
mA mE mA mE mA mE mA mE mA mE
FedAvg (McMahan et al., 2017) 0.780 0.273 0.803 0.273 0.660 0.636 0.617 0.177 0.088 1.91
DisTrans (Yuan et al., 2022) 0.815 - 0.707 - 0.898 - 0.699 - 0.267 -
UEFL (Ours) 0.920 0.149 0.850 0.167 0.942 0.0239 0.720 0.0222 0.326 0.655
@@ 4 Experimental Results
Experimental Setup. As discussed in (Kairouz et al., 2021b; Zhou et al., 2023), there are two predominant forms of data heterogeneity in federated learning: feature heterogeneity and label heterogeneity. Our UEFL focuses on tacking feature heterogeneity, and we mainly discuss feature heterogeneity in this section. The discussion for label heterogeneity and the comparison with VHL (Tang et al., 2022), FedBR (Guo et al., 2023b) are in the Appendix I.
Similar to Rotated MNIST (Ghifary et al., 2015), which creates six domains through counter-clockwise rotations of 0°, 15°, 30°, 45°, 60°, and 75° on MNIST, we employ similar technique to introduce feature heterogeneity on five different datasets: MNIST, FMNIST, CIFAR10, GTSRB, and CIFAR100, to validate our framework’s robustness. In our experiments, we create three domains by counter-clockwise rotating the datasets by 0° ($\mathcal{D}_{1}$), -50° ($\mathcal{D}_{2}$), and 120° ($\mathcal{D}_{3}$). We sampled three data silos from each domain (i.e. totally 9 silos), and data silos for CIFAR100 contain 4000 images each, while the other datasets consist of 2000 images per silo. Besides the regular training with multi-domain data silos, we also test UEFL for domain generalization (DG) task on Rotated MNIST (Ghifary et al., 2015) and PACS (Li et al., 2017) datasets, which contains four distinct domains: art painting (A), cartoon (C), photo (P), and sketch (S).
For RGB datasets like GTSRB, CIFAR10, and CIFAR100, we adopt a pretrained VGG16 model in multi-domain training. In contrast, for grayscale datasets such as MNIST and FMNIST, lacking pretrained models, we design a convolutional network comprising three ResNet blocks, training it from scratch. And for DG, we adopt a pretrained ResNet18 for both datasets. Initial codebook sizes are set to 32 for MNIST and 64 for the remaining datasets, with an equivalent number of codewords added in each subsequent iteration. While additional iterations may converge within 5 rounds, we extend this to 20 for enhanced experimental clarity. The uncertainty evaluation is conducted 20 times using a dropout rate of 0.1, with thresholds $\gamma$ set at 0.3 for MNIST, 0.1 for FMNIST, GTSRB, and CIFAR100, and 0.2 for CIFAR10, to fine-tune performance. These experiments are performed on a machine with two NVIDIA A6000 GPUs.
Evaluation Metrics. We compute the mean Top-1 accuracy (mA), mean entropy (mE), and mean perplexity (mP) across all clients to facilitate a direct comparison, where perplexity quantifies the utility of codewords.
@@ 4.1 Extensible Codebook
Discretization for Heterogeneous FL. To show the effectness of discretization to tackle the data heterogeneity in FL, we design a toy experiment on MNIST. Temporarily setting aside federated learning’s privacy considerations, we directly discretized the features for each client using the distinct codebooks based on its originating domain. With this discretization of VQ-FedAvg, the mean accuracy was improved from 0.834 to 0.907 with the reduction of uncertainty, demonstrating the effectiveness of feature discretization in enhancing performance within a heterogeneous federated learning context, as shown in LABEL:fig:code1.
Table 2: Comparison with different methods for DG. Results are on six domains of Rotated MNIST, four domains of PACS and their average. Our approach is compared with baselines: FedAvg, FedSR, and FedIIR.
Methods Rotated MNIST PACS
$\mathcal{M}_{0}$ $\mathcal{M}_{15}$ $\mathcal{M}_{30}$ $\mathcal{M}_{45}$ $\mathcal{M}_{60}$ $\mathcal{M}_{75}$ Ave. A C P S Ave.
FedAvg (McMahan et al., 2017) 82.7 98.2 99 99.1 98.2 89.9 94.5 78 73 92 79 80.3
FedSR (Nguyen et al., 2022) 84.2 98.0 98.9 99.0 98.3 90.0 94.7 83 75 94 82 83.4
FedIIR (Guo et al., 2023a) 83.8 98.2 99.1 99.1 98.5 90.8 95.0 83 76 94 82 83.7
UEFL (ours) 88.1 97.3 97.6 97.8 97.9 93.2 95.3 81 80 94 82 84.5
Extensible Codebook v.s. Static Large Codebook. To validate our extensible codebook’s superiority over starting with a large codebook, we ensured both methods ended with the same number of codewords through experiments. Results on CIFAR100 in LABEL:fig:code2 demonstrate the difficulties associated with a larger initial codebook in codeword selection for image features. Conversely, gradually expanding the codebook significantly improved codeword differentiation, yielding better outcomes, such as enhanced accuracy (from 0.13 to 0.34), reduced uncertainty (0.78 vs. 1.66 for the static approach), and increased utilization of codewords.
Codebook Initialization. Section 3.1 highlights our UEFL’s capability for efficient codeword initialization via K-means, utilizing features from a finetuned encoder. The efficacy of initialization is validated in LABEL:fig:code3 with results on MNIST, showing enhancements across all metrics.
4.2 UEFL for Multi-Domain Learning
We conducted comparative experiments on five datasets with introduced feature heterogeneity against leading algorithms, specifically the baseline FedAvg (McMahan et al., 2017) and DisTrans (Yuan et al., 2022). For accuracy comparison, DisTrans generally exhibits better performance than FedAvg, making it our primary point of comparison. Regarding uncertainty comparison, because DisTrans lacks Dropout layers, precluding uncertainty evaluation, we exclusively compare uncertainty metrics with FedAvg.
Performance. The results in Table 1 provide a comprehensive comparison, illustrating that UEFL surpasses all other SOTA methods in both accuracy and uncertainty reduction. Specifically, UEFL achieves accuracy improvements ranging from 3% to 22.1% over DisTrans. Our UEFL improves uncertainty compared to FedAvg, achieving reductions by 38.83%-96.24%. LABEL:fig:silo1 and LABEL:fig:silo2 details performance across each data silo, highlighting UEFL’s effectiveness in elevating the accuracy and degrading the uncertainty.
Table 3: UEFL is scalable for 50 clients and beats all SOTA methods: FedAvg, FedSR and FedIIR.
Methods #Clients Backbone Domains Average
$\mathcal{M}_{0}$ $\mathcal{M}_{15}$ $\mathcal{M}_{30}$ $\mathcal{M}_{45}$ $\mathcal{M}_{60}$ $\mathcal{M}_{75}$
FedAvg (McMahan et al., 2017) 50 ResNet18 77.9 95.9 96.9 97 96 81.2 90.8
FedSR (Nguyen et al., 2022) 50 ResNet18 78.3 95.7 96.3 97.1 96 84 91.2
FedIIR (Guo et al., 2023a) 50 ResNet18 84 96.8 97.7 97.7 97.4 84.5 93
UEFL (ours) 50 ResNet18 86.4 95.5 96.4 96.9 94.7 90.6 93.42
Table 4: UEFL is also scalable for 100 clients and beats all SOTA methods following the experimental setup in FedCR.
Methods #Clients EMNIST-L FMNIST CIFAR10 CIFAR100
FedAvg (McMahan et al., 2017) 100 95.89 88.15 76.83 32.08
FedSR (Nguyen et al., 2022) 100 86.22 85.55 61.47 40.82
FedCR (Zhang et al., 2023) 100 97.47 93.78 84.74 62.96
UEFL (ours) 100 98.29 93.93 86.11 63.37
Codewords Perplexity. LABEL:fig:silo3 presents a perplexity comparison between our UEFL and FedAvg, illustrating enhanced codebook utilization after assigning new codewords to $\mathcal{D}_{3}$. This adjustment not only benefits $\mathcal{D}_{3}$ but also improves the codebook utilization for $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$.
Computation Overhead. Our approach introduces only a small codebook, thus incurring negligible memory and computational overheads. Specifically, for the CIFAR10 dataset, the parameter count for the baseline FedAvg model is 14.991M, whereas our UEFL model slightly increases to 15.491M, indicating a tiny memory increment of 3.34%. In terms of runtime, UEFL also exhibits a minimal increase from 16.154ms to 16.733ms (3.58% increase). These findings underscore UEFL’s suitability for deployment on edge devices. More details are included in Appendix G.
4.3 UEFL for Domain Generalization (DG)
For DG task, the trained model needs to be evaluated on an out-of-distribution domain and we follow the evaluation method in (Nguyen et al., 2022; Guo et al., 2023a). Specifically, we perform “leave-one-domain-out” experiments, where we choose one domain as the target domain, train the model on all remaining domains, and evaluate it on the chosen domain. Each source domain is treated as a client.
As shown in Table 2, our UEFL enhanced mean accuracy on the RotatedMNIST dataset, elevating it from 0.945 to 0.953. This performance exceeds that of FedSR (Nguyen et al., 2022) at 0.947 and FedIIR (Guo et al., 2023a) at 0.95. Similarly, on the PACS dataset, UEFL improved mean accuracy from 0.803 to 0.8453, surpassing FedSR’s 0.834 and FedIIR’s 0.837. These results underscore UEFL’s efficacy in tackling feature heterogeneity and superior performance on the federated DG task, beating state-of-the-art methods.
4.4 Ablation Study
Imbalanced Clients. We constructed an experimental setup with three data silos from $\mathcal{D}_{1}$ and one each from $\mathcal{D}_{2}$ and $\mathcal{D}_{3}$, totaling five silos. Our UEFL can also improve both accuracy (from 0.508 to 0.828) and uncertainty (from 0.256 to 0.105) in this scenario. Detailed results are in Appendix K.
Large Number of Clients. We follow (Guo et al., 2023a) to further partition the five training domains of Rotated MNIST into 50 sub-domains, and adopt the setup from (Zhang et al., 2023) to evaluate performance on 100 clients. In both cases, UEFL achieves the highest mean accuracy, outperforming all baselines, as shown in Table 3 and Table 4, demonstrating its scalability to a larger number of clients.
Number of codewords and segments. We investigate the impact of varying the number of initialized codewords in our extensible codebook, to balance accuracy with runtime efficiency in K-means initialization. In LABEL:fig:para1 for GTSRB, initializing with 32 codewords provides comparable accuracy and uncertainty metrics. For more complex datasets, we enhance selection capacity using codeword segmentation. LABEL:fig:para2 demonstrates that segmenting codewords into 4 parts leads to enhanced performance on CIFAR100.
Uncertainty Threshold. In our UEFL, the uncertainty evaluator plays a pivotal role in identifying heterogeneous data without needing direct data access, with the threshold selection being critical. As illustrated in LABEL:fig:para3, a lower threshold imposes stricter criteria, pushing the model to achieve higher performance. However, it’s important to recognize that beyond a certain point, further reducing the threshold may not significantly enhance outcomes but will increase computational overhead. Thus, in such cases, there is a trade-off between runtime and performance.
5 Conclusion
In this work, we address the challenge of data heterogeneity among silos within federated learning setting by introducing an innovative solution: an extensible codebook designed to map distinct data distributions using varied codeword pools. Our proposed framework, Uncertainty-Based Extensible-Codebook Federated Learning (UEFL), leverages this extensible codebook through an iterative process that adeptly identifies data from unknown distributions via uncertainty evaluation and enriches the codebook with newly initialized codewords tailored to these distributions. The iterative nature of UEFL, coupled with efficient codeword initialization using K-means, ensures codewords are closely matched with the actual data distribution, thereby expediting model convergence. This approach allows UEFL to rapidly adjust to new and unseen data distributions, enhancing adaptability. Our comprehensive evaluation across various prominent datasets showcases UEFL’s effectiveness.
Acknowledgements
This work was supported in part by the joint funding program between the National University of Singapore and the University of Toronto. We gratefully acknowledge their support.
Impact Statement
Appendix A Discretization for Generalization
Theoretically, the inclusion of the discretization process offers two key advantages: (1) enhanced noise robustness, and (2) reduced underlying dimensionality. These benefits are demonstrated in the following two theorems. (Liu et al., 2021).
Notation: $\boldsymbol{h}$ is input vector, $\boldsymbol{h}\in\mathcal{H}\in\mathcal{R}^{m}$. $L$ is the size of codebook, $G$ is the number of segments, $q(\cdot)$ is discretization process, $\phi(\cdot)$ is any function (model). Given any family of sets $S=\{S_{1},...,S_{K}\}$ with $S_{1},...,S_{K}\subseteq\mathcal{H}$, we define $\phi_{k}^{S}$ by $\phi_{k}^{S}=\mathbb{1}\{\boldsymbol{h}\in S_{k}\}\phi(\boldsymbol{h})$ for all $k\in[K]$, where $[K]=\{1,...,K\}$. And we denote by $(Q_{k})_{k\in[L^{G}]}$ all the codewords.
Theorem 1: (with discretization) $Let$ $S_{k}=\{Q_{k}\}$ $for$ $all$ $k\in[L^{G}]$. $Then$, $for$ $any$ $\delta>0$, $with$ $probability$ $at$ $least$ $1-\delta$ $𝑜𝑣𝑒𝑟$ $an$ $iid$ $draw$ $of$ $n$ $examples$ $(\boldsymbol{h}_{i})_{i=1}^{n}$, $the$ $following$ $holds$ $for$ $any$ $\phi:\mathcal{R}^{m}\rightarrow\mathcal{R}$ $and$ $all$ $k\in[L^{G}]:$ $if$ $|\phi_{k}^{S}(\boldsymbol{h})|\leq\alpha$ $for$ $all$ $\boldsymbol{h}\in\mathcal{H}$, $then$
[…]
$where$ $no$ $constant$ $is$ $hidden$ $in$ $\mathcal{O}$.
Theorem 2: (without discretization) $Assume$ $that$ $||\boldsymbol{h}||_{2}\leq R_{\mathcal{H}}$ $for$ $all$ $\boldsymbol{h}\in\mathcal{H}\in\mathcal{R}^{m}$. $Fix$ $\mathcal{C}\in argmin_{\overline{\mathcal{C}}}\{|\overline{\mathcal{C}}|:\overline{C}\subseteq\mathcal{R}^{m},\mathcal{H}\subseteq\cup_{c\in\overline{\mathcal{C}}}\mathcal{B}[c]\}$ $where$ $\mathcal{B}[\boldsymbol{c}]=\{\boldsymbol{x}\in\mathcal{R}^{m}:||\boldsymbol{x}-\boldsymbol{c}||_{2}\leq R_{\mathcal{H}}/(2\sqrt{n})\}$. $Let$ $S_{k}=\mathcal{B}[\boldsymbol{c}_{k}]$ $for$ $all$ $k\in[|\mathcal{C}|]$ $where$ $\boldsymbol{c}_{k}\in\mathcal{C}$ $and$ $\cup_{k}\{\boldsymbol{c}_{k}\}=\mathcal{C}$. $Then$, $for$ $any$ $\delta>0$, $with$ $probability$ $at$ $least$ $1-\delta$ $𝑜𝑣𝑒𝑟$ $an$ $iid$ $draw$ $of$ $n$ $examples$ $(\boldsymbol{h}_{i})_{i=1}^{n}$,$the$ $following$ $holds$ $for$ $any$ $\phi:\mathcal{R}^{m}\rightarrow\mathcal{R}$ $and$ $all$ $k\in[|\mathcal{C}|]:$ $if$ $|\phi_{k}^{S}(\boldsymbol{h})|\leq\alpha$ $for$ $all$ $\boldsymbol{h}\in\mathcal{H}$ $and$ $|\phi_{k}^{S}(\boldsymbol{h})-\phi_{k}^{S}(\boldsymbol{h}^{\prime})|\leq\varsigma_{k}||\boldsymbol{h}-\boldsymbol{h}^{\prime}||_{2}$ $for$ $all$ $\boldsymbol{h},\boldsymbol{h}^{\prime}\in S_{k}$, $for$ $all$ $\boldsymbol{h},\boldsymbol{h}^{\prime}\in S_{k}$, $then$
[…]
$where$ $no$ $constant$ $is$ $hidden$ $in$ $\mathcal{O}$ $and$ $\overline{\varsigma_{k}}=\varsigma_{k}(\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}\{h_{i}\in\mathcal{B}[c_{k}]\})$.
Based on these two theorems, we can determine that the performance gap between training and test data is smaller when discretization is applied, due to the following two points:
  • There is an additional error without discretization (i.e. $\frac{\overline{\varsigma_{k}}R_{\mathcal{H}}}{\sqrt{n}}$) in the bound of Theorem 2. This error disappears with discretization in the bound of Theorem 1 as the discretization process reduces the sensitivity to noise.
  • The discretization process reduces the underlying dimensionality of $m\text{ln}(4\sqrt{nm})$ without discretization (in Theorem 2) to that of $G\text{ln}(L)$ with discretization (in Theorem 1). Since the number of discretization heads $G$ (eg. $G$ is 1, 2, or 4 in our case) is always much smaller than the number of dimensions $m$, the inequality $G\text{ln}(L)\leq m\text{ln}(4\sqrt{nm})$ consistently holds.
Appendix B Discretization for non-iid Federated Learning
Notation: Consider a federated setting with $K$ clients. Client $k$ has $n_{k}$ samples independently drawn from its own distribution $P_{k}$. Total samples of all clients $n=\sum_{k=1}^{K}n_{k}$. The global distribution is $\overline{P}=\sum_{k=1}^{K}\frac{n_{k}}{n}P_{k}$.
Theorem 3: (with discretization) $if$ $|\phi_{k}^{S}(\boldsymbol{h})|\leq\alpha$ $for$ $all$ $\boldsymbol{h}\in\mathcal{H}$, $then$, $for$ $any$ $\delta>0$, $with$ $probability$ $at$ $least$ $1-\delta$:
[…]
$where$ $\nu^{(q)}=\frac{1}{K}\sum_{k=1}^{K}\text{Div}(P_{k}^{(q)},\overline{P}^{(q)})$, $denoting$ $the$ $KL$ $divergence$ $between$ $the$ $client$ $distributions$ $and$ $the$ $global$ $distribution$ $after$ $discretization$.
Theorem 4: (without discretization) $Assume$ $that$ $||\boldsymbol{h}||_{2}\leq R_{\mathcal{H}}$, then for any $\delta>0$, with probability $\geq 1-\delta$,
[…]
$where$ $\nu=\frac{1}{K}\sum_{k=1}^{K}\text{Div}(P_{k},\overline{P})$, $denoting$ $the$ $KL$ $divergence$ $between$ $the$ $client$ $distributions$ $and$ $the$ $global$ $distribution$.
The KL divergence term $\nu^{(q)}$ after discretization is significantly smaller than the original $\nu$ due to discretization-induced invariance. Therefore, discretization not only improves robustness to noise and reduces dimensionality, but importantly, it explicitly and effectively mitigates the adverse effects of data heterogeneity typical in non-IID federated learning environments.
Appendix C Different data heterogeneity for UEFL
C.1 UEFL for Multi-Domain Learning
[…]
mA mE mA mE mA mE mA mE mA mE
FedAvg $\mathcal{D}_{1}$ 0.874 0.212 0.801 0.246 0.670 0.623 0.676 0.172 0.110 1.74
All 0.780 0.273 0.803 0.273 0.660 0.636 0.617 0.177 0.088 1.91
DisTrans $\mathcal{D}_{1}$ 0.856 - 0.721 - 0.898 - 0.721 - 0.289 -
$\mathcal{D}_{2}$ 0.799 - 0.705 - 0.900 - 0.719 - 0.261 -
$\mathcal{D}_{3}$ 0.789 - 0.694 - 0.897 - 0.659 - 0.251 -
All 0.815 - 0.707 - 0.898 - 0.699 - 0.267 -
All 0.920 0.149 0.850 0.167 0.942 0.0239 0.720 0.0222 0.326 0.655
C.2 Lower Data heterogeneity on CIFAR100
In Table 5, CIFAR100 exhibits poor performance due to the highly heterogeneous experimental setup. To provide a detailed comparison with the baseline, we used a VGG16 backbone to test CIFAR100 under multiple settings: (1) Local Training: all data trained together without a distributed setting; (2) FedAvg (w/o hete): CIFAR100 split into 5 clients, each with 10,000 images, to evaluate FedAvg performance; (3) FedAvg (w/ hete): images for the 5 clients were rotated by -30°, -15°, 0°, 15°, and 30°, respectively, to introduce data heterogeneity, and FedAvg performance was evaluated; (4) UEFL (w/ hete): tested under the same heterogeneous setup. We trained models from scratch and with pre-trained weights. Results are presented in Table 6, showing that UEFL consistently outperforms FedAvg in both cases.
Table 6: Under lower data heterogeneity on CIFAR100, UEFL continues to outperform FedAvg for both training from scratch and using pre-trained weights.
Training Strategy Local training FedAvg (w/o hete) FedAvg (w/ hete) UEFL (w/ hete)
From scratch 0.3852 0.2447 0.0852 0.1062
Pre-trained 0.6604 0.6496 0.5005 0.5619
C.3 Different data heterogeneity on CIFAR100
By progressively increasing the rotation angles to simulate greater data heterogeneity, we evaluated UEFL’s performance under varying levels of heterogeneity. As shown in Table 7, while overall performance decreases with higher heterogeneity, UEFL consistently outperforms FedAvg, with the performance gap widening as heterogeneity increases, demonstrating UEFL’s superiority in addressing data heterogeneity.
Table 7: With different rotation angles, UEFL keep outperforms FedAvg.
Rotation Angles FedAvg UEFL
{-10°, -5°, 0°, 5°, 10°} 0.5935 0.6232
{-20°, -10°, 0°, 10°, 20°} 0.5469 0.6039
{-30°, -15°, 0°, 15°, 30°} 0.5106 0.56
{-40°, -20°, 0°, 20°, 40°} 0.4799 0.5311
{-50°, -25°, 0°, 25°, 50°} 0.4494 0.5074
{-60°, -30°, 0°, 30°, 60°} 0.4368 0.4939
{-70°, -35°, 0°, 35°, 70°} 0.4188 0.4737
Appendix D Optimal training epochs of baselines
To fully demonstrate the efficacy of our UEFL, besides evaluating baselines with the same total training epochs as UEFL, we remove the additional training epochs from UEFL iterations and obtain the optimal performance, for more fair comparison.
Table 8: UEFL outperforms all baselines without additional training epochs.
Methods MNIST FMNIST GTSRB CIFAR10 CIFAR100
mA mE mA mE mA mE mA mE mA mE
FedAvg 0.782 0.261 0.801 0.289 0.657 0.645 0.618 0.173 0.093 1.74
UEFL (Ours) 0.920 0.149 0.850 0.167 0.942 0.0239 0.720 0.0222 0.326 0.655
Appendix E Deep Ensembles
We evaluated Deep Ensembles (Lakshminarayanan et al., 2017) by creating 5 ensembles to assess uncertainty for our method. Table 9 shows a comparison between Deep Ensembles and Monte Carlo Dropout. From the results, we observe that the accuracy when using Deep Ensembles is quite similar to Monte Carlo Dropout, apart from the stochastic variations. This is expected, as the accuracy is not directly impacted by the choice of uncertainty evaluation method. However, the uncertainty values for Deep Ensembles are higher than those for Monte Carlo Dropout, likely due to the use of only 5 ensembles for evaluation to reduce computational time. In conclusion, while the accuracy is comparable, Deep Ensembles require significantly more computational resources due to the need to train multiple networks. Therefore, Monte Carlo Dropout is a more efficient and suitable choice for our approach.
Table 9: Comparison of Deep Ensembles and Monte Carlo Dropout for uncertainty evaluation.
Methods MNIST FMNIST GTSRB CIFAR10 CIFAR100
mA mE mA mE mA mE mA mE mA mE
Monte Carlo Dropout 0.920 0.149 0.850 0.167 0.942 0.0239 0.720 0.0222 0.326 0.655
Deep Ensemble 0.926 0.211 0.853 0.289 0.940 0.041 0.717 0.043 0.331 0.873
Appendix F Neural Collapse
We conducted experiments on MNIST and CIFAR-100 based on the framework presented in (Papyan et al., 2020). According to the paper, when training continues until the training error reaches 0 (i.e. training accuracy exceeds 99.9% for MNIST/CIFAR-10), the Terminal Phase of Training (TPT) begins, during which neural collapse (NC) emerges. To validate this, we first conducted local training by training on all data together. Our results confirmed the paper’s claim that additional training beyond the zero-error point leads to improved performance. We then extended these experiments to the federated learning setting. The detailed results are presented in Table 10.
From these experiments, we observed the following findings: 1. The dropout layer must be removed; otherwise, the training accuracy cannot exceed 99.9% (e.g., the final training accuracy for MNIST is limited to 96% with dropout). 2. Compared to local training, more training epochs are required to reach the TPT in federated learning. For example, on CIFAR-100 with data heterogeneity, federated learning requires 44 rounds of training (44 × 5 epochs), while local training achieves TPT in 38 epochs. 3. Although neural collapse yields improved performance, UEFL consistently outperforms it, especially on more complex datasets like CIFAR-100. This is partly because removing dropout layers for neural collapse increases the risk of overfitting.
Table 10: Comparison with neural collapse (NC).
Method MNIST CIFAR100
zero-error 0.9375 0.3473
last epoch (NC) 0.9584 0.3482
UEFL (Ours) 0.9778 0.5074
Appendix G Computation Cost
Table 11 compares FedAvg and UEFL. The parameter count for UEFL in this table represents the final model size, including 256 codewords after two iterations starting from 64. In our experiments, 256 is the largest final codebook size across all datasets. For simpler datasets like MNIST, UEFL demonstrates an even lower computation cost.
Table 11: With different rotation angles, UEFL keep outperforms FedAvg.
Method #Params (M) CPU runtime (ms) GPU runtime (ms)
FedAvg 14.991 16.102 16.154
UEFL (Ours) 15.491 16.611 16.733
Appendix H Convergence Curves
For the experiments on the CIFAR10 dataset, in Figure 11, at round 40, after we assign new codewords, rapid performance gains are evident. Remarkably, the training process demonstrates swift convergence, typically within just five rounds. For illustrative clarity and to underscore the differential impact, we extend the training to 20 rounds in subsequent iterations, showcasing the accelerated and effective adaptation of our approach. In addition, after 60 rounds, even if we keep adding new codewords, the increased perplexity denotes a higher utilization of the codebook. However, there is no significant improvement in accuracy or uncertainty. Figure 11 also shows a large boost with our UEFL on MNIST.
Figure 11: Learning curves of MNIST and CIFAR10. (a) Accuracy of MNIST. (b) Accuracy of CIFAR10. (c) Uncertainty of CIFAR10. (d) Perplexity of CIFAR10.
For the experiments on the FMNIST and GTSRB datasets, as depicted in Figure 12, we introduce new codewords to the codebook only once. Notably, there is a clear ”performance jump” evident in all six figures, showcasing the rapid adaptation of our UEFL to new data distributions.
Figure 12: Learning curves of FMNIST and GTSRB. (left) Accuracy results. (mid) Uncertainty results. (right) codeword loss results.
Appendix I Our UEFL for Label Heterogeneity
Similar to (Tang et al., 2022; Guo et al., 2023b), we introduce label heterogeneity with dirichlet distribution ($\alpha=0.1$). The results in Table 12 show that our UEFL can also tackle the label heterogeneity when compared to FedAvg and performs better than VHL for CIFAR10 even if it cannot perform as well as FedBR.
Table 12: Comparison with different methods on data with label heterogeneity.
Method FMNIST CIFAR10
FedAvg VHL UEFL (ours) FedAvg VHL FedBR UEFL (ours)
mA 87.45 91.52 90.59 58.99 61.23 64.61 62.67
Appendix J K-means Initialization
Figure 13 illustrates this concept: gray points represent features from the trained encoder, clustered according to their data distributions. While direct data access is restricted, differentiation by uncertainty allows us to identify and utilize the centroids of these clusters via K-means for codeword initialization.
And to bolster the robustness of our methodology, we dissect features into smaller segments—using factors like 2 or 4—to pair them with multiple codewords, thus covering the entirety of a feature vector as illustrated in Figure 3. This segmentation exponentially increases the codeword pool to $n^{2}$ or $n^{4}$, ensuring a robust representation capacity.
Appendix K Imbalanced Clients
As shown in Table 13, our UEFL also works for imbalanced data silos when there are three clients sampled from the same domain. Both accuracy and uncertainty get improved, especially for the third domain.
Table 13: Our UEFL improves the performance of unbalanced data.
Data Silo FedAvg UEFL
Acc Entropy Acc Entropy
$\mathcal{D}_{1}$ $s_{1_{a}}$ 0.964 0.0312 0.952 0.0291
$s_{1_{b}}$ 0.936 0.0252 0.974 0.0170
$s_{1_{c}}$ 0.964 0.0499 0.944 0.0308
$\mathcal{D}_{2}$ $s_{2_{a}}$ 0.796 0.1477 0.836 0.1261
$\mathcal{D}_{3}$ $s_{3_{a}}$ 0.508 0.2560 0.828 0.1048
Appendix L UEFL Optimization